OpenStudy (anonymous):

What is the common difference of the following arithmetic sequence? -4, -5.6, -7.2, -8.8,…

OpenStudy (texaschic101):

-1.6

OpenStudy (anonymous):

thnks

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

What are the next 3 terms of the following sequence? 3, -1, -5, -9,…

OpenStudy (anonymous):

What is the common difference of the following arithmetic sequence? 5, 4.5, 4, 3.5,…

OpenStudy (texaschic101):

-4 is common difference, so next three terms are : -4 - 9 = - 13 -4 - 13 = - 17 - 4 - 17 = -21 common difference is -0.5

OpenStudy (anonymous):

thnks so much... now what about this one What are the next 3 terms of the following sequence? 3, -1, -5, -9,…

OpenStudy (texaschic101):

already did it......next terms are -13, -17, -21

OpenStudy (anonymous):

o ok..... didnt see it, thnks so much!

OpenStudy (texaschic101):

anytime

OpenStudy (anonymous):

do you think you can help me with more

OpenStudy (texaschic101):

possibly

OpenStudy (anonymous):

Find the three arithmetic means between -5 and -25.

OpenStudy (anonymous):

Find S20 for the series -1 + -3 + -5 + -7 +…

OpenStudy (texaschic101):

-5 + 4d = -25 4d = -25 + 5 4d = -20 d = -5 -5, (-10, -15, -20), -25 <-- 3 terms are in parenthesis ================================== an = a1 + (n - 1) d n = 20 a1 = -1 d = -2 a20 = -1 + (20 - 1) * -2 a20 = -1 + (19) * -2 a20 = -1 - 38 a20 = -39 <-- this is just the 20th term...not the sum Be patient...I am getting there sn = n(a1 + an)/2 n = 20 a1 = -1 s20 = 20(-1 + (-39) / 2 s20 = 20 (-40) / 2 s20 = - 800/2 s20 = -400 <-- this is the sum

OpenStudy (anonymous):

thks.. you are great! ima give you more

OpenStudy (anonymous):

Find the 6th term of the sequence with t1 = -4 and tn = 5tn-1.

OpenStudy (anonymous):

Find S8 for the series -2 + -10 + -50 + -250 +…